### v2

@@ 8.2 Absolutely continuous solutions to the Fokker-Planck equation
Let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$. A (regular enough) function $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is said to solve the non-homogeneous Fokker-Planck Equation with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, if
@@
For every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and every $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, there exists a unique mapping $p\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{n},\mathbb{R}))$ which solves FPEf with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, and with $\frac{\partial p}{\partial t}\in L^{2}(0,T;H^{-1}(\mathbb{R}^{n},\mathbb{R}))$. In addition, the following first parabolic estimate holds:
$\|p(t,\cdot)\|^{2}_{L^{2}}+\int^{t}_{0}$ $\|p(t,\cdot)\|^{2}_{H^{1}}\;\mathrm{d}t\leq$
(27) $\leq C\big{(}\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{L^{2}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T],$
where $C\big{(}\|(a,b)\|_{\mathcal{H}_{m}}\big{)}>0$ is a constant which continuously depends on $\|(a,b)\|_{\mathcal{H}_{m}}$. Finally, if $f=0$ and $\bar{p}$ is a non-negative density in $L^{2}(\mathbb{R}^{n},\mathbb{R})$, then
@@
where the measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ solves SDE for some coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, and the mapping $f:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is regular enough. In the case the coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ of SDE are to be learned, one does not have perfect knowledge of (28), and therefore the error between (28) and its counterpart in which $(X_{x})_{x\in\mathbb{R}^{n}}$ is replaced with the solution to SDE stemming from rather learned coefficients must be estimated. This gap is filled with the following:
@@
For any coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, let $X^{a,b}:\mathbb{R}^{n}\times\Omega\to S$ denote the unique measurable mapping which solves SDE with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ (see Theorem 8.6). There exists an operator
$\mathcal{O}:$ $\mathcal{H}^{+}_{m}\to L^{\infty}(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R})^{*})\cong L^{\infty}(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))$
@@
In addition, if $p_{a,b}\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))$ denotes the unique solution to FDE0 with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ and $\bar{p}=p_{0}$ (which uniquely exists thanks to Theorem 8.9), for every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ the mapping
@@
is measurable and in $L^{1}([0,T],\mathbb{R})$, and it satisfies, for every $(a_{1},b_{1}),(a_{2},b_{2})\in\mathcal{H}^{+}_{m}$,
@@
Fix any $(a,b)\in\mathcal{H}^{+}_{m}$. For $t\in[0,T]$ and $\varphi\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, we define
@@ 8.3 Additional regularity of solutions to the Fokker-Planck equation
In Section 5, we dealt with solutions $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ of FPE0 which enjoy higher regularity properties, and specifically for which $p\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$, and for which FPE0 holds pointwise. Below, we are going to show such additional attributes stem from the properties of coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ listed in Theorem 2.1.
First, for every $(a,b)\in\mathcal{H}^{+}_{m}$ and almost every $t\in[0,T]$ we introduce the notation:
$(\mathcal{L}^{a,b}_{t})^{*}u(t,y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}\frac{\partial^{2}\big{(}a^{\alpha}_{ij}u\big{)}}{\partial y_{j}\partial y_{i}}(t,y)-\sum^{n}_{i=1}\frac{\partial\big{(}b_{i}u\big{)}}{\partial y_{i}}(t,y),\quad u\in L^{2}(0,T;H^{2}(\mathbb{R}^{n},\mathbb{R})),$
@@
$\hskip 131.31903pt\textnormal{and}\quad\frac{\partial^{2}\big{(}a^{\alpha}_{ij}u\big{)}}{\partial y_{r}\partial y_{s}}=a^{\alpha}_{ij}\frac{\partial^{2}u}{\partial y_{r}\partial y_{s}}+\frac{\partial a^{\alpha}_{ij}}{\partial y_{s}}\frac{\partial u}{\partial y_{r}}$
$\hskip 161.45782pt+\frac{\partial a^{\alpha}_{ij}}{\partial y_{r}}\frac{\partial u}{\partial y_{s}}+u\frac{\partial^{2}a^{\alpha}_{ij}}{\partial y_{r}\partial y_{s}}\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R}),$
@@
Under the setting and notation of Theorem 8.9, if furthermore $\bar{p}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$, then the unique solution $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ to FPE0 with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$ is additionally such that
@@
In particular, the mapping $p$ additionally satisfies the following Strong Fokker-Planck Equations with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$:
@@ A.2 Proofs of Section 8
The proof is standard and based on the classical Lions scheme (see, e.g., [25, 10, 13]), therefore we mainly focus on deriving the constant $C\big{(}\|(a,b)\|_{\mathcal{H}_{m}}\big{)}=C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}$ in (8.9) and the properties of $p$ when $f=0$ and $\bar{p}$ is a non-negative density in $L^{2}(\mathbb{R}^{n},\mathbb{R})$. For this, consider the Gelfand triple
@@
$(t;u,v)\mapsto\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a^{\alpha}_{ij}(t,y)\frac{\partial u}{\partial y_{j}}(y)\frac{\partial v}{\partial y_{i}}(y)\;\mathrm{d}y$
$\hskip 86.11084pt+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial v}{\partial y_{i}}(y)u(y)\;\mathrm{d}y.$
@@
$a^{\alpha}_{ij}(t,\cdot)p(t,\cdot)\in H^{1}(\mathbb{R}^{n},\mathbb{R}),\quad\textnormal{with}\quad\frac{\partial(a^{\alpha}_{ij}p)}{\partial y_{j}}(t,\cdot)=p(t,\cdot)\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,\cdot)+a^{\alpha}_{ij}(t,\cdot)\frac{\partial p}{\partial y_{j}}(t,\cdot),$
@@
$\int_{\mathbb{R}^{n}}a^{\alpha}_{ij}(t,y)p(t,y)$ $\frac{\partial^{2}\varphi}{\partial y_{j}\partial y_{i}}(y)\;\mathrm{d}y=$
$=-\int_{\mathbb{R}^{n}}\left(a^{\alpha}_{ij}(t,y)\frac{\partial p}{\partial y_{j}}(t,y)+p(t,y)\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)\right)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y,$
showing that $p$ solves FPEf with coefficients $(a,b)\in\mathcal{H}^{+}_{m}$, $\alpha>0$.
@@
$=-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a^{\alpha}_{ij}(t,y)\frac{\partial(p^{+}-p^{-})}{\partial y_{j}}(t,y)\frac{\partial(-p^{-})}{\partial y_{i}}(t,y)\;\mathrm{d}y$
$\hskip 64.58313pt+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial(p^{+}-p^{-})}{\partial y_{i}}(t,y)p(t,y)^{-}\;\mathrm{d}y$
@@
$=\underset{k\to\infty}{\lim}\ \bigg{(}\int_{\mathbb{R}^{n}}\bar{p}(y)\varphi_{k}(y)\;\mathrm{d}y+\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a^{\alpha}_{ij}(t,y)\frac{\partial^{2}\varphi_{k}}{\partial y_{i}\partial y_{j}}(y)p(t,y)\;\mathrm{d}y$
@@
(36) $\leq\|f(t,\cdot)\|_{L^{2}}\left\|\frac{\partial p_{k}}{\partial t}(t,\cdot)\right\|_{L^{2}}-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a^{\alpha}_{ij}(t,y)\frac{\partial p_{k}}{\partial y_{j}}(t,y)\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(t,y)\;\mathrm{d}y$
$\quad-\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial}{\partial y_{i}}\left(\frac{\partial p_{k}}{\partial t}\right)(t,y)\ p_{k}(t,y)\;\mathrm{d}y.$
@@
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}a^{\alpha}_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)\;\mathrm{d}y\;\mathrm{d}s=$
$=-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}\bigg{(}\frac{\mathrm{d}}{\mathrm{d}s}\left(a^{\alpha}_{ij}\frac{\partial p_{k}}{\partial y_{j}}\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)-\dot{a}^{\alpha}_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\partial p_{k}}{\partial y_{i}}(s,y)$
$\hskip 161.45782pt-a^{\alpha}_{ij}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{j}}\right)(s,y)\ \frac{\partial p_{k}}{\partial y_{i}}(s,y)\bigg{)}\;\mathrm{d}y\;\mathrm{d}s,$
and since $a^{\alpha}_{ji}=a^{\alpha}_{ij}$ for every $i,j=1,\dots,n$, we may compute
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}a^{\alpha}_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)\;\mathrm{d}y\;\mathrm{d}s\leq$
$\leq-\sum^{n}_{i,j=1}\frac{1}{2}\int_{\mathbb{R}^{n}}\left(a^{\alpha}_{ij}(t,y)\frac{\partial p_{k}}{\partial y_{j}}(t,y)\frac{\partial p_{k}}{\partial y_{i}}(t,y)-a^{\alpha}_{ij}(0,y)\frac{\partial p_{k}}{\partial y_{j}}(0,y)\frac{\partial p_{k}}{\partial y_{i}}(0,y)\right)\;\mathrm{d}y\;\mathrm{d}s$
@@
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}$ $a^{\alpha}_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}(s,y)\right)\;\mathrm{d}y\;\mathrm{d}s\leq$
(37) $\leq-\frac{\alpha}{2}\|\nabla p_{k}(t,\cdot)\|^{2}_{L^{2}}+C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),$
where $C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}>0$ is a constant which continuously depends on $\alpha$ and $\|(a,b)\|_{\mathcal{H}_{m}}$ uniquely. Below, we will implicitly overload the constant $C\big{(}\alpha,\|(a,b)\|_{\mathcal{H}_{m}}\big{)}$.
@@
$)\displaystyle-\int_{\mathbb{R}^{n}}\bigg{(}\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)$$\bigg{)}\frac{\partial\varphi}{\partial y_{i}}(y)p_{k}(t,y)\;\mathrm{d}y=$
$=\int_{\mathbb{R}^{n}}\frac{\partial}{\partial y_{i}}\left(\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a^{\alpha}_{ij}}{\partial y_{j}}-b_{i}\right)p_{k}\right)(t,y)\ \varphi(y)\;\mathrm{d}y.$

### v3

@@ 8.2 Absolutely continuous solutions to the Fokker-Planck equation
Let $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$. A (regular enough) function $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is said to solve the non-homogeneous Fokker-Planck Equation with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, if
@@
For every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ and every $\bar{p}\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, there exists a unique mapping $p\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))\cap L^{2}(0,T;H^{1}(\mathbb{R}^{n},\mathbb{R}))$ which solves FPEf with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, and with $\frac{\partial p}{\partial t}\in L^{2}(0,T;H^{-1}(\mathbb{R}^{n},\mathbb{R}))$. In addition, the following first parabolic estimate holds:
$\|p(t,\cdot)\|^{2}_{L^{2}}+\int^{t}_{0}$ $\|p(s,\cdot)\|^{2}_{H^{1}}\;\mathrm{d}s\leq$
(29) $\leq C\big{(}{\color[rgb]{0,0,0}\alpha},\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{L^{2}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),\quad t\in[0,T],$
with $C\big{(}{\color[rgb]{0,0,0}\alpha},\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}\big{)}>0$ continuously depending on ${\color[rgb]{0,0,0}\alpha}$ and $\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}$. Finally, if $f=0$ and $\bar{p}$ is a non-negative density in $L^{2}(\mathbb{R}^{n},\mathbb{R})$, then
@@
where the measurable mapping $X:\mathbb{R}^{n}\times\Omega\to S$ solves SDE for some coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, and the mapping $f:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ is regular enough. In the case the coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ of SDE are to be learned, one does not have perfect knowledge of (30), and therefore the error between (30) and its counterpart in which $(X_{x})_{x\in\mathbb{R}^{n}}$ is replaced with the solution to SDE stemming from rather learned coefficients must be estimated. This gap is filled with the following:
@@
For any coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, let $X^{a,b}:\mathbb{R}^{n}\times\Omega\to S$ denote the unique measurable mapping which solves SDE with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ (see Theorem 8.6). There exists an operator
$\mathcal{O}:$ ${\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}\to L^{\infty}(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R})^{*})\cong L^{\infty}(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))$
@@
In addition, if $p_{a,b}\in C(0,T;L^{2}(\mathbb{R}^{n},\mathbb{R}))$ denotes the unique solution to FDE0 with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ and $\bar{p}=p_{0}$ (which uniquely exists thanks to Theorem 8.9), for every $f\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R})$ the mapping
@@
is measurable and in $L^{1}([0,T],\mathbb{R})$, and it satisfies, for every $(a_{1},b_{1}),(a_{2},b_{2})\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$,
@@
Fix any $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$. For $t\in[0,T]$ and $\varphi\in L^{2}(\mathbb{R}^{n},\mathbb{R})$, we define
@@ 8.3 Additional regularity of solutions to the Fokker-Planck equation
In Section 5, we dealt with solutions $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ of FPE0 which enjoy higher regularity properties, and specifically for which $p\in H^{m+1,2(m+1)}([0,T]\times\mathbb{R}^{n},\mathbb{R})$, and for which FPE0 holds pointwise. Below, we are going to show such additional attributes stem from the properties of coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ listed in Theorem 2.1.
First, for every $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ and almost every $t\in[0,T]$ we introduce the notation:
$(\mathcal{L}^{a,b}_{t})^{*}u(t,y)\triangleq\frac{1}{2}\sum^{n}_{i,j=1}\frac{\partial^{2}\big{(}a_{ij}u\big{)}}{\partial y_{j}\partial y_{i}}(t,y)-\sum^{n}_{i=1}\frac{\partial\big{(}b_{i}u\big{)}}{\partial y_{i}}(t,y),\quad u\in L^{2}(0,T;H^{2}(\mathbb{R}^{n},\mathbb{R})),$
@@
$\hskip 131.31903pt\textnormal{and}\quad\frac{\partial^{2}\big{(}a_{ij}u\big{)}}{\partial y_{r}\partial y_{s}}=a_{ij}\frac{\partial^{2}u}{\partial y_{r}\partial y_{s}}+\frac{\partial a_{ij}}{\partial y_{s}}\frac{\partial u}{\partial y_{r}}$
$\hskip 161.45782pt+\frac{\partial a_{ij}}{\partial y_{r}}\frac{\partial u}{\partial y_{s}}+u\frac{\partial^{2}a_{ij}}{\partial y_{r}\partial y_{s}}\in L^{2}([0,T]\times\mathbb{R}^{n},\mathbb{R}),$
@@
Under the setting and notation of Theorem 8.9, if furthermore $\bar{p}\in H^{2m+1}(\mathbb{R}^{n},\mathbb{R})$, then the unique solution $p:[0,T]\times\mathbb{R}^{n}\to\mathbb{R}$ to FPE0 with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$ is additionally such that
@@
In particular, the mapping $p$ additionally satisfies the following Strong Fokker-Planck Equations with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$:
@@ A.2 Proofs of Section 8
The proof is standard and based on the classical Lions scheme (see, e.g., [25, 10, 13]), therefore we mainly focus on deriving the constant ${\color[rgb]{0,0,0}C\big{(}\alpha,\|(a-\alpha I,b)\|_{\mathcal{H}_{m}}\big{)}}$ in (8.9) and the properties of $p$ when $f=0$ and $\bar{p}$ is a non-negative density in $L^{2}(\mathbb{R}^{n},\mathbb{R})$. For this, consider the Gelfand triple
@@
$(t;u,v)\mapsto\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a_{ij}(t,y)\frac{\partial u}{\partial y_{j}}(y)\frac{\partial v}{\partial y_{i}}(y)\;\mathrm{d}y$
$\hskip 86.11084pt+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial v}{\partial y_{i}}(y)u(y)\;\mathrm{d}y.$
@@
$a_{ij}(t,\cdot)p(t,\cdot)\in H^{1}(\mathbb{R}^{n},\mathbb{R}),\quad\textnormal{with}\quad\frac{\partial(a_{ij}p)}{\partial y_{j}}(t,\cdot)=p(t,\cdot)\frac{\partial a_{ij}}{\partial y_{j}}(t,\cdot)+a_{ij}(t,\cdot)\frac{\partial p}{\partial y_{j}}(t,\cdot),$
@@
$\int_{\mathbb{R}^{n}}a_{ij}(t,y)p(t,y)$ $\frac{\partial^{2}\varphi}{\partial y_{j}\partial y_{i}}(y)\;\mathrm{d}y=$
$=-\int_{\mathbb{R}^{n}}\left(a_{ij}(t,y)\frac{\partial p}{\partial y_{j}}(t,y)+p(t,y)\frac{\partial a_{ij}}{\partial y_{j}}(t,y)\right)\frac{\partial\varphi}{\partial y_{i}}(y)\;\mathrm{d}y,$
showing that $p$ solves FPEf with coefficients $(a,b)\in{\color[rgb]{0,0,0}\mathcal{H}^{+,\alpha}_{m}}$, $\alpha>0$.
@@
$=-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a_{ij}(t,y)\frac{\partial(p^{+}-p^{-})}{\partial y_{j}}(t,y)\frac{\partial(-p^{-})}{\partial y_{i}}(t,y)\;\mathrm{d}y$
$\hskip 64.58313pt+\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial(p^{+}-p^{-})}{\partial y_{i}}(t,y)p(t,y)^{-}\;\mathrm{d}y$
@@
$=\underset{k\to\infty}{\lim}\ \bigg{(}\int_{\mathbb{R}^{n}}\bar{p}(y)\varphi_{k}(y)\;\mathrm{d}y+\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a_{ij}(t,y)\frac{\partial^{2}\varphi_{k}}{\partial y_{i}\partial y_{j}}(y)p(t,y)\;\mathrm{d}y$
@@
(38) $\leq\|f(t,\cdot)\|_{L^{2}}\left\|\frac{\partial p_{k}}{\partial t}(t,\cdot)\right\|_{L^{2}}-\sum^{n}_{i,j=1}\int_{\mathbb{R}^{n}}\frac{1}{2}a_{ij}(t,y)\frac{\partial p_{k}}{\partial y_{j}}(t,y)\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(t,y)\;\mathrm{d}y$
$\quad-\sum^{n}_{i=1}\int_{\mathbb{R}^{n}}\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)\right)\frac{\partial}{\partial y_{i}}\left(\frac{\partial p_{k}}{\partial t}\right)(t,y)\ p_{k}(t,y)\;\mathrm{d}y.$
@@
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}a_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)\;\mathrm{d}y\;\mathrm{d}s=$
$=-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}\frac{1}{2}\bigg{(}\frac{\mathrm{d}}{\mathrm{d}s}\left(a_{ij}\frac{\partial p_{k}}{\partial y_{j}}\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)-\dot{a}_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\partial p_{k}}{\partial y_{i}}(s,y)$
$\hskip 161.45782pt-a_{ij}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{j}}\right)(s,y)\ \frac{\partial p_{k}}{\partial y_{i}}(s,y)\bigg{)}\;\mathrm{d}y\;\mathrm{d}s,$
and since $a_{ji}=a_{ij}$ for every $i,j=1,\dots,n$, we may compute
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}\sum^{n}_{i,j=1}a_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}\right)(s,y)\;\mathrm{d}y\;\mathrm{d}s\leq$
$\leq-\sum^{n}_{i,j=1}\frac{1}{2}\int_{\mathbb{R}^{n}}\left(a_{ij}(t,y)\frac{\partial p_{k}}{\partial y_{j}}(t,y)\frac{\partial p_{k}}{\partial y_{i}}(t,y)-a_{ij}(0,y)\frac{\partial p_{k}}{\partial y_{j}}(0,y)\frac{\partial p_{k}}{\partial y_{i}}(0,y)\right)\;\mathrm{d}y\;\mathrm{d}s$
@@
$-\int^{t}_{0}\int_{\mathbb{R}^{n}}$ $\sum^{n}_{i,j=1}a_{ij}(s,y)\frac{\partial p_{k}}{\partial y_{j}}(s,y)\frac{\mathrm{d}}{\mathrm{d}s}\left(\frac{\partial p_{k}}{\partial y_{i}}(s,y)\right)\;\mathrm{d}y\;\mathrm{d}s\leq$
(39) $\leq-\frac{\alpha}{2}\|\nabla p_{k}(t,\cdot)\|^{2}_{L^{2}}+C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}\left(\|\bar{p}\|^{2}_{H^{1}}+\int^{t}_{0}\|f(s,\cdot)\|^{2}_{L^{2}}\;\mathrm{d}s\right),$
where $C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}>0$ continuously depends on $\alpha$ and $\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}$ uniquely. Below, we will implicitly overload the constant $C\big{(}\alpha,\|({\color[rgb]{0,0,0}a-\alpha I},b)\|_{\mathcal{H}_{m}}\big{)}$.
@@
$)\displaystyle-\int_{\mathbb{R}^{n}}\bigg{(}\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}(t,y)-b_{i}(t,y)$$\bigg{)}\frac{\partial\varphi}{\partial y_{i}}(y)p_{k}(t,y)\;\mathrm{d}y=$
$=\int_{\mathbb{R}^{n}}\frac{\partial}{\partial y_{i}}\left(\left(\sum^{n}_{j=1}\frac{1}{2}\frac{\partial a_{ij}}{\partial y_{j}}-b_{i}\right)p_{k}\right)(t,y)\ \varphi(y)\;\mathrm{d}y.$